Adger Corporation is a service company that measures its output based on the number of customers served. The company provided the following fixed and variable cost estimates that it uses for budgeting purposes and the actual results for May as shown below:

Fixed Element
per Month
Variable Element per Customer Served Actual Total
for May
Revenue $ 6,100 $ 223,500
Employee salaries and wages $ 68,000 $ 1,500 $ 126,000
Travel expenses $ 600 $ 20,400
Other expenses $ 47,000 $ 44,300

When preparing its planning budget the company estimated that it would serve 35 customers per month; however, during May the company actually served 40 customers.

3.

value:
4.00 points

Required information

3. What amount of travel expenses would be included in Adger’s flexible budget for May?

5. What net operating income would appear in Adger’s flexible budget for May?

7. What is Adger’s employee salaries and wages spending variance for May? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

12. What amount of travel expenses would be included in Adger’s planning budget for May?

Travel Expenses

Variable element per customer served

Planned activity

Amount in planning Budget
Answer #1

3) Amount of travel expense in flexible budget = 40*600 = 24000

5) Net income under flexible budget = (4000*40-68000-47000) = 45000

7) Employee salaries and wages spending variance = (40*1500+68000)-126000 = 2000 F

12) Travel expense =

variable element per customer served = 600

Planned activity = 35

Amount in planning budget = 21000